#1 Each basket will contain the following items:

DVD $8.00

Pasta $2.00

Sauces $3.00

Cheeses $4.00

Pasta tongs $5.00

Wines $8.00

Baskets $4.00

Cellophane $1.00

Ribbons $1.00

1 gal. of gas $4.00

I can make 20 gift baskets. I estimate my fixed costs to include the salary I want to achieve $200 and the utilities that I may use to produce these which is around $25.  

If I want to realize a 25% return on sales, where should I price my product for the holiday season?

Answering the question: What percent of each dollar sales = profit? A 20% return on sales means that for every $1.00 of sales you created, 20 cents of it was profit.  

#2 For the holiday season, I plan to make and sell a game. Each game will need:

Piping $15.00

Golf balls $6.00

Cords $2.00

The plan is to sell each of these games for $38. I can make 40 games before the holidays. My only fixed costs will be a salary I hope to make of $200.00. Calculate my return on sales on this effort: Using the same formula, but you are solving for the actual percentage return on sales.

?% = Total Revenue – Total Cost

                         Total Revenue
Answer #1
Variable Cost $
DVD 8
Pasta 2
Sauces 3
Cheeses 4
Pasta tongs 5
Wines 8
Baskets 4
Cellophane 1
Ribbons 1
Gas 4
Total Variable Cost per Basket (A) 40
Number of baskets(B) 20
Total Variable cost(C)=(A)*(B) 800
Fixed Cost
Salary 200
Utility 25
Total Fixed Cost(D) 225
Total Cost(C)+(D) 1025

Return on sales=25%

Let sale be 100

Then Profit=25

Cost=100-25=75

It means cost vaule is 75

In our case cost is 1025 that is equals to 75%

Sale=1025/75%=1366.67

Price of the product=Sale/Units

=1366.67/20

=$68.33 per Basket

#2

Variable Cost $
Piping 15
Golf balls 6
Cords 2
Variable Cost per basket(A) 23
number of games(B) 40
Total Variable Cost (C)=(A)*(B) 920
Total Fixed Cost i. e. Salary (D) 200
Total Cost (C)+(D) 1120
Sale Revenue=38*40 1520

Return on Sales=(Total Revenue-Total Cost)/Total Revenue

=(1520-1120)/1520

=26.32%

"Please Rate if it helps you or let me know if anything else"